OpenStudy (anonymous):

Will medal and Fan! What is the slope of the line passing through the points (1, –5) and (4, 1)? A.2 B.-4/5 C.5/4 D.-2 Thanks :D

OpenStudy (anonymous):

slope formula\[=\frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (anonymous):

\[Let~say(1,-5)=(x_1,y_1)\]\[Let~say(4,1)=(x_2,y_2)\]substitute this value into the formula to find the slope

OpenStudy (anonymous):

\[=\frac{ 1-(-5) }{ 4-1 }\]

OpenStudy (anonymous):

\[=\frac{ 1+5 }{ 4-1 }\]\[=\frac{ 6 }{ 3 }\]simplify it @Xsciencegirlx